A 670m long eureka wire has resistance of 250 ohms. What length of the same wire will have a resistance of 110 ohms?


1
Expert's answer
2022-03-14T13:06:34-0400

Answer

Resistance is given by

"R=\frac{\rho l}{A}"

So in above case length of wire is given

"l'=\frac{R'l}{R}\\=\frac{110*670}{250}\\=294.8m"